OpenStudy (anonymous):

In right triangle JKL, m < J = 90° and tan K = sqrt 5/4 Find the length of the hypotenuse. √21 3 16√5

OpenStudy (anonymous):

@KamiBug

OpenStudy (anonymous):

@StudyGurl14

OpenStudy (anonymous):

@thomaster

OpenStudy (studygurl14):

 Do you know what the tan ratio is in terms of which sides?

OpenStudy (anonymous):

opp and adj @StudyGurl14

OpenStudy (studygurl14):

Correct.

OpenStudy (anonymous):

got it already can you help me with couple other questions?

OpenStudy (anonymous):

thnx btw

OpenStudy (studygurl14):

So if... \(\large\tan\ K=\Large\frac{\sf opposite}{\sf adjacent}\) And... \(\large\tan\ K=\Large\frac{\sqrt{5}}{4}\) What are the lengths?

OpenStudy (studygurl14):

Nice job. :)

OpenStudy (studygurl14):

Yeah, tag me to a new thread for the other question